OpenStudy (anonymous):

I need someone to help me finish my thought process. I am graphing system of inequalities. x≥30,y≥20 3x+5y≤340

OpenStudy (anonymous):

What do i do next?

OpenStudy (anonymous):

Can i simplify? Or do I just graph?

OpenStudy (anonymous):

great, thanks!

OpenStudy (turingtest):

you can graph x≥30,y≥20 right away, just graph x=30 and y=20 and shade above it because it is 'greater-than'

OpenStudy (anonymous):

x=30 would be parallel to y axis and y=30 would be parallel to x axis.

OpenStudy (turingtest):

the last one will be helpful to put in slope-intercept form\[y\le-\frac35x+68\]

OpenStudy (anonymous):

Thank you both for your help. I think I understand it now! :D

OpenStudy (turingtest):

*could be helpful

OpenStudy (anonymous):

For, \[3x+5y≤340 \implies \large \frac{x}{\frac{340}3} + \frac{y}{\frac{340}5} \le 1\] now, this is very easy to plot.

OpenStudy (anonymous):

THANKS!

OpenStudy (anonymous):

Glad to help :)